OpenStudy (anonymous):

Please help! When a certain number n is divided by 3, the result is the same as when n is decreased by 3 .what is n ?

OpenStudy (helder_edwin):

\[ \large n\div3=n-3 \]

OpenStudy (anonymous):

Help me please

Directrix (directrix):

@Jessie123456789 Okay, I'll help.

OpenStudy (anonymous):

Thanks

Directrix (directrix):

How do we write these words in math symbols --> When a certain number n is divided by 3.

Directrix (directrix):

Be brave. Take a guess.

OpenStudy (anonymous):

So what n and can u explain it

OpenStudy (anonymous):

N divided by 3 = n - 3

OpenStudy (anonymous):

I think

Directrix (directrix):

n is a variable and represents the number we are looking for. So, I am asking you what n divided by 3 looks like. Say, 5 divided by 3. It looks like 5/3 in symbols. Note: I cannot write out the solution without asking questions or I will be banished from this site. 0.o So, what does n divided by 3 look like?

OpenStudy (anonymous):

N/3

OpenStudy (anonymous):

Is it right ?

Directrix (directrix):

Yes, it is right. Good. N divided by 3 = N - 3 ---> correct except we to get rid of the word divide and replace it by the symbol for division. So, N/3 = N - 3 We have to solve this for N

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

How do we do that?

Directrix (directrix):

We need to get the N terms together on one side of the equation so ... N/3 = N - 3 Subtract N from both sides: N/3 - N = -3 What is (1/3)N - 1N equal to? Just like 1/3 - 1 but with an N.

OpenStudy (anonymous):

I don't know

Directrix (directrix):

Okay. We'll multiply through by 3 and avoid the hassle of fractions. N/3 - N = -3 3( N/3 - N) = 3(-3) N - 3N = -9 Your turn: What is N - 3N ? (We're almost finished)

OpenStudy (anonymous):

Um I think it is -12?

OpenStudy (anonymous):

Or maybe -9

Directrix (directrix):

1 - 3 = ? One minus three = ?

OpenStudy (anonymous):

Oh 2

OpenStudy (anonymous):

I mean -22

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Not -22 that was a mistake

Directrix (directrix):

3-1 = 2; 1-3 = -2. -----------------

OpenStudy (anonymous):

So n= -2?

Directrix (directrix):

N - 3N = -9 -2N = -9 N = -9/(-2) N = ?

OpenStudy (anonymous):

-2 ?

Directrix (directrix):

-9 divided by -2 = 9/2. N = 9/2 I'll put together all the parts of the work we did so that you can study it later.

OpenStudy (anonymous):

Oh okay thanks you help me so much !

Directrix (directrix):

N/3 = N - 3 3( N/3 - N) = 3(-3) N - 3N = -9 -2N = -9 N = -9/(-2) N = 9/2 --> Answer

Directrix (directrix):

There's all our work in one place. Glad to help.

OpenStudy (anonymous):

My big sister didn't even want to help and she's in 8 th grade and I'm in 7th

Directrix (directrix):

There will come a time when she will need your help on something. You can remind her of this. c:

OpenStudy (anonymous):

I know thanks again

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

But I will still help her even though she is so mean I am not gonna be like her at all trust me

Directrix (directrix):

@Jessie123456789 The Middle School years are weird sometimes. Just be who you are. And, yes, help those who ask. Great outlook on life, I say.